OpenStudy (anonymous):

A toy company has determined that the revenue generated by a particular toy is modeled by the following equation: r(X)=11x-0.025x^2 The variable x is measured in thousands of toys produced, and r(x) is measured in thousands of dollars. What is the maximum revenue the company can earn with this toy? Give your answer in dollars, without using commas.

OpenStudy (anonymous):

r '(x) = 11 - 0.05(x) = 0 x = 11(20) = 220 r(220) = 11(220) - 0.025(220)^2 = $1210000